It has been a long time for me.
Looking for X in this equation.
134000/X = 1576470
please show the steps. i know the answer. I am justing trying to see the steps.

Answer by Alan3354(69443)

134000/X = 1576470
Divide by 10
13400/X = 157647
Multiply by x
13400 = 157647x
x = 13400/157647
x =~ 0.085